OpenStudy (anonymous):

what is the solution to this system of linear equations? y=-9x+20 y=6x-25

OpenStudy (anonymous):

6x-25=20-9x 15x=45 x=3 put this value in any equation y=6*3-25=-7

OpenStudy (dape):

We know that y should equal both right hand sides, so equate them: \[ -9x+20=6x-25 \\ 45=15x \\ x=3 \] Now we can get y by inserting x into one of the original equations: \[ y=6x-25=3*6-25=18-25=-7 \]